OpenStudy (anonymous):

find the y intercept and the x intercept for -3x+y=5... I know this is an easy one, but I am having problems

OpenStudy (anonymous):

soz forgot the minus sign lets do it again

OpenStudy (anonymous):

\[(y=5-3x) = (y=-3x+5)\] so the intercept is 5 because y=mx+c where c is the intercept

OpenStudy (anonymous):

To find x intercept set y=0 so \[0=-3x+5 \]\[-5=-3x\]\[x=\frac{-5}{-3}\] so when y =0 x=-5\-3 or 5\3